OpenStudy (anonymous):

What equation is solved by the graphed systems of equations? the graph is at (-1,-4)

OpenStudy (anonymous):

the answer choices are 7x + 3 = x + 3 7x - 3 = x - 3 7x + 3 = x - 3 7x - 3 = x + 3

OpenStudy (anonymous):

are any of those variables supposed to be y? or is there a graph to look at or anything? because that is confusing. sorry.